OpenStudy (twizttiez):

Variable y varies directly with x^2, and y = 48 when x = 4. Which graph represents the quadratic variation?

OpenStudy (twizttiez):

Phew thanks for coming i really need help @geerky42

geerky42 (geerky42):

"Variable y varies directly with x^2, and y = 48 when x = 4" So you are given \(y = kx^2\), where k is unknown. Firstly you need to solve for k. Plug in y=48 and x=4. \((48) = k(4)^2\) Can you do that?

OpenStudy (twizttiez):

3?

OpenStudy (twizttiez):

@geerky42

OpenStudy (twizttiez):

@campbell_st

OpenStudy (campbell_st):

you have calculated k correctly so the equation is \[y = 3x^2\]

OpenStudy (campbell_st):

I haven't looked at the choices... but using the equation you can find some points it passes through x: -2 : -1: 0 : 1 : 2 -------------------- y then look at the choices and see if the parabola passes through them hope it helps

OpenStudy (twizttiez):

So my answer is C then? @campbell_st

OpenStudy (campbell_st):

I just viewed and it does seem to match the table values

OpenStudy (twizttiez):

Ok thanks so much sorry that i didn't respond sooner os is running really slow